OpenStudy (tomy12345):

ind the product of (x − 3)2. (1 point) x2 + 6x + 9 x2 − 9 x2 − 6x + 9 x2 + 9

OpenStudy (tomy12345):

@Qwertty123 @AloneS

alones (alones):

(x−3)(2) =(x+−3)(2) =(x)(2)+(−3)(2) =2x−6 Wait or do you need to graph it?

OpenStudy (tomy12345):

thanks